# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v1

@@ 3.2 Dynamic Masked Degradation
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity helps to mitigate noise memorization.
In this task, one of the major challenges is semantic pollution, where undesired visual semantics could be perceived by the T2I model and thus “pollute” the personalized concept. As shown in Figure 2(a.i), the target concept (i.e., person) could be severely disturbed by the owner of the target component (i.e., eye), resulting in a hybrid person. Unfortunately, directly masking out the regions other than the target concept and component would damage the overall visual context, thus leading to overfitting and weird compositions in Figure 2(a.ii). In light of that, the undesired visual semantics of reference images should be processed properly. Hence, we propose Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics (see Figure 3), aiming at suppressing the T2I model’s perception for them while maintaining overall visual contexts (see Figure 2(a.iii)).
Degradation Imposition. In each training step, DM-Deg imposes degradation in the regions outside segmentation masks for each reference image. There are various types of degradation that can be adopted to perturb the visual semantics of an image, such as noise, blur, and geometric distortions, but not all of them are easy to use and compatible with mask operations. In DM-Deg, we choose to employ Gaussian noise due to its simplicity. For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$. Note that here the pixel values of $I_{nk}$ range from $-1$ to $1$. Then, the imposition of degradation is conducted as
@@
where $\odot$ indicates element-wise multiplication and $\alpha_{d}\in[0,1]$ is a dynamic weight used to regulate the degradation intensity for $I_{nk}$. In this way, we can obtain a randomly degraded image $\hat{I}_{nk}$ where the original visual contexts are generally retained. Encountering with $\hat{I}_{nk}$, it is more difficult for the T2I model to perceive undesired visual semantics in out-of-mask regions, since these semantics would be randomly perturbed with Gaussian noise at each training step.
Dynamic Intensity. Unfortunately, the T2I model may gradually memorize the introduced noise when learning meaningful visual semantics, thus making noise appear in generated images (see Figure 4(a)). Such behavior also aligns with previous observations on deep networks (Arpit et al., 2017). Hence, we design a descending scheme to dynamically regulate the intensity of the imposed noise during training. This scheme adopts an exponential curve that maintains a relatively large intensity in the early steps and decreases dramatically in the later ones. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as
@@
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ is a factor to regulate the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$ tuned within the powers of $2$. Using such a scheme of dynamic intensity, we can effectively prevent semantic pollution and meanwhile alleviate the memorization of the introduced noise, achieving better generation performance (see Figure 4(b)).
Figure 5: Visualization of the learning process. (a) The vanilla learning paradigm lapses into overemphasizing the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
@@ 3.3 Dual-Stream Balancing
Another primary challenge in this task is semantic imbalance, which arises from the inherent visual semantic disparity between the target concept and component. A concept generally has richer visual semantics than a component (e.g. person vs. hair). while the semantic richness of a component might be greater than a concept in some cases (e.g., simple tower vs. intricate roof). This imbalance complicates the joint learning process which could overemphasize on either concept or component, leading to incoherent generation of the concept or component (see Figure 5(a)). To address this challenge, we design Dual-Stream Balancing (DS-Bal), which establishes a dual-stream learning paradigm with online and momentum denoising U-Nets (see Figure 3) to balance visual semantic learning of the concept and component, improving personalization fidelity (see Figure 5(b)).
@@
Sample-wise Min-Max Optimization. From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ for all the samples. Unfortunately, this indiscriminate optimization does not allocate adequate learning efforts to the sample that is more challenging to learn, gradually leading to an imbalanced learning process. To overcome this issue, DS-Bal utilizes the online denoising U-Net to learn only the visual semantics of the hardest-to-learn sample at each training step. Inheriting the weights of the original denoising U-Net warmed up by joint learning, here the online denoising U-Net $\epsilon_{\theta}$ only optimizes the maximum masked diffusion loss among $N$ samples, which is defined as
$\mathcal{L}_{\text{diff-max}}=\max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (6)
@@
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization (Razaviyayn et al., 2020). The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
@@
Selective Preserving Regularization. At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which can be expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as
@@
$\mathcal{L}_{\text{pres}}=\mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (7)
@@
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA (Tarvainen & Valpola, 2017) with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other sample while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Finally, using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as
@@
4 Experiments
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and the compared methods for various domains. MagicTailor generally achieves promising text alignment, strong identity fidelity, and high generation quality. More results are provided in Appendix D.
@@ 4.1 Experimental Setup
Dataset, Implementation, and Evaluation. For a systemic investigation, we collect a dataset from various domains, including characters, animation, buildings, objects, and animals. We employ Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model. Reference images are resized to 512 $\times$ 512, and the LoRA rank and alpha are set to 32. For the warm-up and DS-Bal stage, we set the training steps to 200 and 300 and the learning rate to 1e-4 and 1e-5, using AdamW (Loshchilov & Hutter, 2017) as the optimizer. To generate evaluation images, we carefully design 20 text prompts covering extensive situations. For each method, we generate 14,720 images to conduct a comprehensive evaluation. To ensure fairness, all the seeds are fixed during training and inference.
Compared Methods. We compare MagicTailor with SOTA methods of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024). For a fair and meaningful comparison, they are adapted to our task with minimal modification, i.e., incorporating the masked diffusion loss (Equation 1) into them. Except for method-specific configurations, all methods use the same implementation above to ensure fairness. Due to the space limit, more details of the experimental setup are provided in Appendix A.
@@ 4.2 Qualitative Comparisons
The qualitative results are presented in Figure 6. It shows that TI, CD, and CLiC mainly suffer from semantic pollution, where undesired visual semantics severely influence the personalized concept. Besides, we can observe that DB and BAS also underperform in this challenging task. These methods exhibit an overemphasis on the concept or component due to semantic imbalance, which even leads to an absence of the target component. Moreover, there is also an interesting observation that imbalanced learning could aggravate the effect of semantic pollution, making the color and texture of the target concept or component mistakenly transferred into an unexpected part of generated images. Compared with these methods, MagicTailor can achieve superior performance in generating text-aligned images that also faithfully reflect the target concept and component, demonstrating its remarkable performance in this newly formulated task.
Table 1: Quantitative comparisons. We compare our MagicTailor with SOTA methods of personalization based on automatic metrics and user study. The best results are marked in bold.
Methods Automatic Metrics User Study
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$ Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
Textual Inversion (Gal et al., 2022) 0.236 0.742 0.620 0.558 5.8% 2.5% 5.2%
DreamBooth (Ruiz et al., 2023) 0.266 0.841 0.798 0.323 15.3% 14.7% 12.5%
Custom Diffusion (Kumari et al., 2023) 0.251 0.797 0.750 0.407 7.1% 7.7% 9.8%
Break-A-Scene (Avrahami et al., 2023) 0.259 0.840 0.780 0.338 10.8% 12.1% 22.8%
CLiC (Safaee et al., 2024) 0.263 0.764 0.663 0.499 4.5% 5.1% 6.2%
MagicTailor (Ours) 0.270 0.854 0.813 0.279 56.5% 57.9% 43.4%
Figure 7: Ablation of loss weights. We report CLIP-T for text alignment, and DreamSim for identity fidelity as it is most similar to human judgments (Fu et al., 2023). For reference, we also present the results of the second-best method in Table 1, highlighting our robustness on loss weights.
4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)). To precisely measure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept (see the detailed setup in Appendix A). As we can see, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment. It should be credited to the utilization of an effective framework tailored to this special task.
User Study. We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality (see the detailed setup in Appendix A). Finally, we collect a total of 3,180 valid answers and report the selected rates in Table 1. It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
@@ 4.3 Quantitative Comparisons
Table 3: Ablation of DS-Bal. We compare DS-Bal with its variants, showing its excellence.
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
@@ 4.4 Ablation Studies
We conduct comprehensive ablation studies of MagicTailor, aiming to verify the capability of the overall pipeline. For more ablation studies on other aspects, please refer to Appendix C.
Figure 8: (a) Decoupled generation. MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components. MagicTailor shows the potential to handle more than one component, highlighting its effectiveness.
@@
Effectiveness of Key Techniques. In Table 4, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Section 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates the significance of these two key techniques. Qualitative results can also refer to Figure 2.
@@
Dynamic Intensity Matters. In Table 4, we explore DM-Deg by comparing it with 1) the mask-out strategy; 2) the fixed intensity; 3) the linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) the dynamic intensity with different $\gamma$. First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution. Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
@@
Momentum Denoising U-Net as a Good Regularizer. In Table 4, we study DS-Bal by modifying the U-Net for regularization as 1) the fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) the fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) the momentum U-Net with other $\beta$. The results demonstrate that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
@@
Sensitivity Analysis of Loss Weights. In Figure 7, we analyze the sensitivity of loss weights in Equation 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on loss weights.
4.5 Further Applications

### v2

@@ 3.2 Dynamic Masked Degradation
Semantic pollution is a major challenge for component-controllable personalization. As shown in Fig. 2 (a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person. Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig. 2 (a.ii)). To address this, undesired visual semantics in reference images must be handled appropriately. We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig. 2 (a.iii)&(c)).
Degradation Imposition.
In each training step, DM-Deg imposes degradation in the out-of-mask region for each reference image. We use Gaussian noise for degradation due to its simplicity. For a reference image $I_{nk}$, we randomly sample a Gaussian noise matrix $G_{nk}\sim\mathcal{N}(0,1)$ with the same shape as $I_{nk}$, where the pixel values of $I_{nk}$ range from $-1$ to $1$. The degradation is then applied as follows:
@@
where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity. While previous works [48, 21] have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context. By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
Dynamic Intensity.
Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig. 4 (a)). This behavior is consistent with previous observations on deep networks [1]. To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training. This scheme follows an exponential curve, maintaining a relatively high intensity in the early stages and decreasing it sharply in later stages. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as:
@@
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ controls the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$. This dynamic intensity scheme effectively prevents semantic pollution and mitigates the memorization of introduced noise, leading to improved generation performance (Fig. 4 (b)&(c)).
@@ 3.3 Dual-Stream Balancing
Another challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component. Concepts generally have richer visual semantics than components (e.g. person vs. hair), but in some cases, the component may have more complex semantics (e.g., simple tower vs. intricate roof). This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig. 5 (a)). To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm with online and momentum denoising U-Nets (Fig. 3) for balanced semantic learning, improving personalization fidelity (Fig. 5 (b) & Fig. 2 (c)).
Figure 5: Learning process visualization. (a) The vanilla learning paradigm tends to overemphasize the easier one. (b) DS-Bal effectively balances the learning of the concept and component.
Sample-Wise Min-Max Optimization.
@@
From a loss perspective, the visual semantics of the concept and component are learned by optimizing the masked diffusion loss $\mathcal{L}_{\text{diff}}$ across all the samples. However, this indiscriminate optimization fails to allocate sufficient learning effort to a more challenging sample, leading to an imbalanced learning process. To address this, DS-Bal uses the online denoising U-Net to focus on learning the hardest-to-learn sample at each training step. Inheriting the weights of the original denoising U-Net, which is warmed up through joint learning, the online denoising U-Net $\epsilon_{\theta}$ optimizes only the sample with the highest masked diffusion loss as:
$\mathcal{L}_{\text{diff-max}}\ =\ \max_{n}\mathbb{E}_{k,\epsilon,t}\Big{[}\big{\|}$ $\epsilon_{n}\odot M_{nk}^{\prime}-$
$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (6)
@@
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization [32]. The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
Selective Preserving Regularization.
@@
At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
@@
$\mathcal{L}_{\text{pres}}\ =\ \mathbb{E}_{n\in S,k,t}\Big{[}\big{\|}\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}-$
$\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})$ $\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (7)
@@
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA [42] with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Overall, DS-Bal can be considered a mechanism to adaptively assign target labels $\epsilon_{n}$ or preserving labels $\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ to different samples, enabling dynamic loss supervision (Fig. 3). Using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as:
@@
4 Experimental Results
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and others across various domains. MagicTailor achieves strong text alignment, identity fidelity, and high-quality generation. More results are in the supplementary material.
@@ 4.1 Experimental Setup
Dataset, Implementation, and Evaluation.
For a systematic investigation, we compile a dataset from diverse domains, including characters, animation, buildings, objects, and animals. We use Stable Diffusion 2.1 [35] as the pretrained T2I model and resize all reference images to $512\times 512$. For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively. Leveraging LoRA fine-tuning, each concept-component pair requires only about five minutes of training on an A100 GPU. For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method. To ensure fairness, all random seeds are fixed during both training and inference.
Compared Methods.
We compare our MagicTailor with state-of-the-art (SOTA) personalization methods, including Textual Inversion (TI) [11], DreamBooth (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements. For a fair comparison, we adapt them to our task with minimal modifications, specifically by incorporating the masked diffusion loss (Eq. 1). Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency. Due to space constraints, additional details on the experimental setup can be found in the supplementary material.
@@ 4.2 Qualitative Comparisons
The qualitative results are shown in in Fig. 6. As observed, TI, CD, and CLiC primarily suffer from semantic pollution, where undesired visual semantics significantly distort the personalized concept. Besides, DB and BAS also struggle in this challenging task, with an overemphasis on either the concept or the component due to semantic imbalance, sometimes even causing the target component to be completely absent. An interesting finding is that imbalanced learning can exacerbate semantic pollution, leading to the color and texture of the target concept or component being mistakenly transferred to unintended parts of the generated images. In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component. To further validate its performance, we provide additional comparisons and a generalizability analysis in the supplementary material.
Table 1: Quantitative comparisons on automatic metrics. MagicTailor can achieve SOTA performance on all four automatic metrics. The best results are marked in bold.
Methods CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ DreamSim $\downarrow$
Textual Inversion [11] 0.236 0.742 0.620 0.558
DreamBooth [36] 0.266 0.841 0.798 0.323
Custom Diffusion [19] 0.251 0.797 0.750 0.407
Break-A-Scene [2] 0.259 0.840 0.780 0.338
CLiC [38] 0.263 0.764 0.663 0.499
MagicTailor (Ours) 0.270 0.854 0.813 0.279
Table 2: Quantitative comparisons on the user study. MagicTailor can also outperform other methods across all three aspects in human evaluation.
Methods Text Align. $\uparrow$ Id. Fidelity $\uparrow$ Gen. Quality $\uparrow$
Textual Inversion [11] 5.8% 2.5% 5.2%
DreamBooth [36] 15.3% 14.7% 12.5%
Custom Diffusion [19] 7.1% 7.7% 9.8%
Break-A-Scene [2] 10.8% 12.1% 22.8%
CLiC [38] 4.5% 5.1% 6.2%
MagicTailor (Ours) 56.5% 57.9% 43.4%
@@
4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely meafsure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept. As we can see in Tab. 1, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment. It should be credited to the utilization of an effective framework tailored to this special task.
User Study.
We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality. Finally, we collect a total of 3,180 valid answers and report the selected rates in Tab. 2. It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
@@ 4.4 Ablation Studies
We conduct comprehensive ablation studies for MagicTailor to verify the capability of the overall pipeline. More ablation studies are included in the supplementary material.
Effectiveness of Key Techniques.
@@
In Tab. 3, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Sec. 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance. Qualitative results can also refer to Fig. 2.
Figure 7: Applications of MagicTailor. (a) Decoupled generation. MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components. MagicTailor shows the potential to handle more than one component, highlighting its effectiveness. (c) Enhancing other generative tools. MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
Dynamic Intensity Matters.
@@
In Tab. 4, we explore DM-Deg by comparing it with 1) mask-out strategy; 2) fixed intensity; 3) linear intensity ($\alpha$ goes from 1 to 0, or from 0 to 1); and 4) dynamic intensity with different $\gamma$. First, the terrible performance of the mask-out strategy verifies that it is not a valid solution for semantic pollution. Notably, the descent linear intensity shows better identity fidelity than its ascent counterpart, which aligns with and validates our observations on noise memorization. Moreover, the dynamic intensity generally shows better results, and it can achieve better overall performance with a proper $\gamma$.
Momentum Denoising U-Net as a Good Regularizer.
@@
In Tab. 5, we study DS-Bal by modifying the U-Net for regularization as 1) fixed U-Net with $\beta=0$ (i.e., the one just after warm-up); 2) fixed U-Net with $\beta=1$ (i.e., the one from the last step); and 3) momentum U-Net with other $\beta$. The results show that employing the U-Net with a high momentum rate can yield better regularization to tackle semantic imbalance, thus leading to excellent performance.
Sensitivity Analysis of Loss Weights.
@@
In Fig. 8, we analyze the sensitivity of loss weights in Eq. 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on these crucial hyperparameters.
Table 5: Ablation of DS-Bal. We compare DS-Bal with potential variants, showing its excellence.
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
4.5 Applications
Decoupled Generation.